OpenStudy (anonymous):

FlybynightSavings.com is offering a savings account that pays 29% compounded continuously. How much interest would a deposit of $4,000 earn over 5 years? (Round your answer to the nearest cent.)

OpenStudy (anonymous):

Compound interest formula: A=P(1+r/n)^nt

OpenStudy (anonymous):

The continuous compound interest formula is A=Pe^rt

OpenStudy (anonymous):

So since this is continuous so we would use that formula, A=4000e^(.29)(5) A=4000e^(1.45) A=17052.46